OpenStudy (anonymous):

if y varies directly as x and y=12 when x=20 then what is y when x is 15

OpenStudy (praxer):

since y is proportional to x so, y= kx where k is proportionality constant. => k= y/x since y/x is a constant term so, 12/20=k also y/15 is also = k => y/15=12/20 =>y=12*15/20 =>y=9 Did you understand it @tevin1 ......:)

OpenStudy (anonymous):

yea thanks

OpenStudy (praxer):

Welcome...:)